Please remember to round your answer to two decimal places.

Problem 9-5 Cost of Equity: Dividend Growth

Summerdahl Resort's common stock is currently trading at $33.00 a share. The stock is expected to pay a dividend of $1.75 a share at the end of the year (D1 = $1.75), and the dividend is expected to grow at a constant rate of 5% a year. What is the cost of common equity? Round your answer to two decimal places.________%

Problem 9-8 WACC

David Ortiz Motors has a target capital structure of 40% debt and 60% equity. The yield to maturity on the company's outstanding bonds is 12%, and the company's tax rate is 40%. Ortiz's CFO has calculated the company's WACC as 11.29%. What is the company's cost of equity capital? Round your answer to two decimal places.______ %

Problem 10-01 NPV

A project has an initial cost of $45,125, expected net cash inflows of $10,000 per year for 9 years, and a cost of capital of 8%. What is the project's NPV? (Hint: Begin by constructing a time line.) Do not round your intermediate calculations. Round your answer to the nearest cent. $______

Problem 10-02 IRR

A project has an initial cost of $60,000, expected net cash inflows of $14,000 per year for 9 years, and a cost of capital of 12%. What is the project's IRR? Round your answer to two decimal places.________ %

Problem 10-03 MIRR

A project has an initial cost of $43,425, expected net cash inflows of $12,000 per year for 10 years, and a cost of capital of 8%. What is the project's MIRR? Do not round intermediate calculations. Round your answer to two decimal places.______%

Problem 10-04 Profitability Index

A project has an initial cost of $61,000, expected net cash inflows of $12,000 per year for 11 years, and a cost of capital of 9%. What is the project's PI? Do not round your intermediate calculations. Round your answer to two decimal places. ________

Problem 10-05 Payback

A project has an initial cost of $59,050, expected net cash inflows of $13,000 per year for 8 years, and a cost of capital of 13%. What is the project's payback period? Round your answer to two decimal places._______ years

Problem 10-06 Discounted Payback

A project has an initial cost of $60,000, expected net cash inflows of $14,000 per year for 7 years, and a cost of capital of 13%. What is the project's discounted payback period? Round your answer to two decimal places.________ years

Please remember to round your answer to two decimal places.
Answer #1

As per rules I am answering the first 4 subparts of the question

9-5: Price = D1/(k-g)

33= 1.75/(k-0.05)

K = 1.75/33 + 0.05 = 10.3%

9-8

WACC = weight of debt *Cost of debt*(1-Tax) + weight of equity*Cost of equity

0.1129 = 0.4*0.12*(1-0.4) + 0.6* Cost of equity

0.1129= 0.0288+ 0.6*Cost of equity

Cost of equity = 14.02%

10-01

NPV =C0+ CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

= C0+ Annual cash flow*(1-1/(1+r)^n)/r

= -45125+ 10000*(1-1/1.08^9)/0.08

=17343.88

10-02

IRR as per excel = 18.12%

AutoSave Book1 Excel (Product Activation Failed) Sign in File Insert Page Layout Formulas Data Review View Help Tell me what you want to do 수 Share Home Calibri Percentage 田FF Conditional Fornat as Cell Insert Delete Fornat FormattingTable Styles- 、. Sort & Find & Paste . 녀 . 의 _ . 로三들經垣臣 Merge & Center-5 . % , : Filter-Select- Clipboard Alignment Number Styles Cells Editing C13 Year 0 CF 14000 14000 14000 14000 14000 14000 14000 14000 14000 4 10 12 13 14 IRR(C2:C11 16 17 Sheet1 Ready + 100 cENG 12:20 AM